One antifreeze solution is 10% alcohol. Another antifreeze solution is 18% alcohol. How many liters of each antifreeze solution should be combined to create 20 liters of antifreeze solution that is 15% alcohol?
Answer by Alan3354(69443)

One antifreeze solution is 10% alcohol. Another antifreeze solution is 18% alcohol. How many liters of each antifreeze solution should be combined to create 20 liters of antifreeze solution that is 15% alcohol?
---------------
t = 10% mix
e = 18% mix
e + t = 20 (20 liters total)
0.1t + 0.18e = 0.15*20 = 3
t + 1.8e = 30
Sub (20-e) for t
20 - e + 1.8e = 30
0.8e = 10
e = 12.5 liters of 18%
t = 7.5 liter of 10%
------------------------
As a check, it "looks right." If you mix it 50-50, it would be 14%, the average of 10% and 18%. Using more of the "stronger" solution makes it more than 14%.